# Agentless [Uncaptioned image] : Demystifying LLM-based Software Engineering Agents

### v1

@@ 2.2 Repair [Uncaptioned image]
In the repair stage, the goal is to produce the correct patch to solve the issue. Following existing work on LLM-based program repair Kolak et al. [2022]; Prenner et al. [2022]; Jiang et al. [2023], we first utilize the identified edit locations and construct a context window of code snippets to provide to the LLM for repair. For example, if the identified location was a class from line 40 to 78, we would produce a context window of [40 - x, 78 + x] where x denotes the context window size. The intuition behind adding the additional code before and after the identified location is to provide the LLM with relevant contextual information for better program repair Xia & Zhang [2022]. If multiple edit locations are identified, we would concatenate these context windows together separated with ``...'' to indicate missing context in the middle.
@@
Patch format. Using the code snippets, we then ask the LLM to generate patches to solve the issue. However, instead of directly producing the entire code snippet to replace the entire given context, Agentless asks the LLM to generate a Search/Replace edit Gauthier [2024]: a simple diff format to efficiently create each patch. Figure 3 shows an example of the Search/Replace format containing two main parts: 1) search: the original code snippet we want to replace and 2) replace: the replacement code snippet we want to replace with. To apply the generated Search/Replace diff to the original file, we can simply match the search code snippet and replace it with the replacement. This simple diff format avoids generating the complete code and instead focuses on producing small edits, which are not only more cost-efficient, but also more reliable and accurate (less chances for hallucination).
Filtering and patch selection. For each issue, Agentless uses the LLM to generate multiple potential patches (starting with greedy and then sample multiple patches with higher temperature). We also apply traditional software engineering technique of regression testing Wong et al. [1997] to run the existing tests in the repository on all the generated patches. Any patches which failed the existing tests can be filtered out as they incorrectly change the correct behavior of previous code. Note, our implementation of this regression test filtering step follows prior work also evaluated on the same benchmark Gauthier [2024]; Chen et al. [2024]. For the remaining patches, Agentless applies a re-ranking approach using majority voting: We first normalize each patch to ignore surface-level differences (e.g., extra spaces, newlines, and comments), and then select the patch with the highest number of occurrences as the final patch for submission. More specifically, to standardize the patch, we begin by parsing both the old and new code (after applying the patch) into abstract syntax trees. Next, we unparse the trees into a canonical source code format with docstrings removed. Finally, we compute the textual diff between the standardized old and new code to get the normalized patch.
@@
Agentless solves repository-level issues using a simple step-by-step procedure. We note here that none of the techniques used by Agentless in isolation are revolutionary, but instead Agentless smartly combines existing techniques to construct an easy-to-understand approach. Different from prior autonomous agent-based tools that involve complex interactions with the environment, Agentless uses a simple two phase approach to first localize and then repair the bug without relying on any agents for decision-making. By conducting localization in a hierarchical manner, Agentless can efficiently and effectively compute the fine-grained locations for editing. Agentless then performs repair by sampling multiple patches using a simple diff format. We filter out any patches with syntax and regression tests errors, and finally select the patch for submission using classic majority voting.
@@ 3 Experimental Setup
Datasets. We evaluate Agentless and baselines using the popular SWE-bench dataset to test the ability to solve real-world software engineering issues. Each problem in SWE-bench requires submitting a patch to solve the underlying issue described in the input issue description. In particular, we focus on the filtered subset SWE-bench Lite, containing 300 problems with tests to evaluate the functional correctness of submitted patch. Furthermore, we also conduct a detailed study (Section 5.1) on the SWE-bench Lite benchmark to not only demonstrate potential issues and biases but also produce a more rigorous filtered set of problems for better evaluation.
@@
Implementation. We implement Agentless using GPT-4o (gpt-4o-2024-05-13OpenAI [2024]. By default, we query the LLM with greedy decoding. During sampling, we use a sampling temperature of $0.8$. For each issue, we first localize to the top three suspicious files, and then localize to an unrestricted number of suspicious classes and functions within these files, all using greedy decoding. Next, to maximize the chances of finding the correct edit locations, we draw four samples of edit locations per issue (i.e., the third step in the localization phase), and combine two sampling runs together to provide more context for repair. This gives us two separate sets of edit locations per issue. For each set, we adopt a context window of $\pm$ 10 lines around each edit location, and generate 21 patches (1 greedy and 20 samples). This results in a total of 42111the answer to the ultimate question of life, the universe, and everything Adams [1995] patches per bug. We adopt the same Search/Replace edit format from prior work Gauthier [2024], and use the built-in Python ast library ast [2024] to perform parsing in our patch normalization step. Due to issues with the original SWE-bench evaluation script at the time of writing, we adopt the SWE-bench-docker Örwall [2024] evaluation setup used by prior tools Gauthier [2024].
@@
Baselines. We compare Agentless against 13 agent-based approaches. These baseline tools represent the state-of-the-art performance on SWE-bench. We include state-of-the-art open-source as well as commercial or closed-source baselines (indicated via a [Uncaptioned image] ). We note here that the majority of the closed-source baselines do not provide any trajectories, just the submission patches. Therefore, we cannot verify the steps taken to arrive at the final patches. Moreover, we also include a simple agentless baseline using retrieval-augmented generation (RAG) proposed as part of SWE-bench Jimenez et al. [2024a] for comparison. In this case, the agentless baseline uses the LLM to directly generate a patch file by providing it with the file content of the most relevant files, retrieved using BM25 Robertson et al. [2009]. Additionally, we also list the underlying LLM used by each tool whenever possible.
@@
Metrics. Following prior work Zhang et al. [2024], we report 1) % Resolved: the percentage of resolved problems in the benchmark, 2) Avg. \$ Cost: average inference cost of running the tool, and 3) Avg. # Tokens: average number of input and output tokens used to query to LLM. Additionally, we also report the % Correct Location: the percent of problems where the patch produced by the tool matches with the edit location of the ground truth developer patch. We compute this metric over three granularities: file, function, and line. We report that a patch contains the correct location if it edits a superset of all locations in the ground truth patch. For baseline tools, we directly use the reported results either from the official leaderboard Jimenez et al. [2024b] or from the tool's official paper/repository.
@@ 4 Evaluation
Table 1: Results on SWE-bench Lite. Note: [Uncaptioned image] indicates approaches that are closed-source (i.e., source code is not released). '-' indicates that the relevant information to compute this has not been released.
Tool LLM % Resolved Avg. \$ Cost Avg. # Tokens % Correct Location
@@
Alibaba Lingma Agent lin [2024] [Uncaptioned image] [Uncaptioned image] GPT-4o+ [Uncaptioned image] Claude-3.5 99 (33.00%) - - 40.0% 58.7% 75.0%
@@
AutoCodeRover-v2 aut [2024] [Uncaptioned image] [Uncaptioned image] GPT-4o 92 (30.67%) - - 35.0% 52.3% 69.3%
@@
Bytedance MarsCode mar [2024] [Uncaptioned image] [Uncaptioned image] GPT-4o 76 (25.33%) - - 37.3% 52.7% 73.7%
@@
Aider Gauthier [2024] [Uncaptioned image] GPT-4o+ [Uncaptioned image] Claude-3 79 (26.33%) - - 35.3% 50.0% 69.7%
AutoCodeRover Zhang et al. [2024] [Uncaptioned image] GPT-4 57 (19.00%) \$0.45 38,663 29.0% 42.3% 62.3%
Claude-3
35 (11.67%) \$3.42 221,258 26.3% 36.0% 48.0%
@@
SWE-agent Yang et al. [2024] [Uncaptioned image] GPT-4o 54 (17.00%) - - - - -
[Uncaptioned image] GPT-4 50 (16.67%) - - 29.0% 39.0% 55.3%
OpenDevin ope [2024b] [Uncaptioned image] GPT-4o 52 (17.33%) - - 27.3% 39.3% 56.7%
Claude-3
@@
RAG Jimenez et al. [2024a] [Uncaptioned image] ChatGPT 1 (0.33%) - - 6.3% 11.3% 27.3%
Agentless [Uncaptioned image] [Uncaptioned image] GPT-4o 82 (27.33%) \$0.34 42,376 34.3% 51.0% 68.7%
@@
Repair performance. Table 1 shows the main evaluation result of Agentless and prior agent-based approaches on SWE-bench Lite. We observe that Agentless is able to solve 82 out of 300 problems (27.33%). While this is not the highest percentage of problems solved on SWE-bench Lite, Agentless is extremely competitive compared with prior agent-based approaches while using a much simpler design and overall technique. It is important to note here that many of the top techniques are closed-source/commercial and did not release any source code to reproduce experiments or even trajectories for further verification. Compared with open-source approaches, Agentless is able to achieve the highest performance of 27.33% (82 / 300) on SWE-bench Lite. Additionally, Agentless only costs on average \$0.34, which is drastically less than prior agent-based approaches. Comparing against the RAG agentless baselines, we see that while Agentless costs slightly more, Agentless is also able to fix way more issues.
Figure 4: Venn diagram for issue fixes
@@
Unique issues fixed. Figure 4 shows the unique issues solved by Agentless compared with the top-performing closed-source/commercial and open-source approaches (``Others'' in Figure 4 indicates all other approaches within each category). First, we see that compared to the open-source agent-based techniques, Agentless is able to fix 15 issues that no other existing open-source agent can resolve, showing the success of using a simple agentless approach in solving difficult issues. Furthermore, even when compared with high-performing commercial approaches, Agentless is still able to offer unique fixes, with even more unique patches than Alibaba Lingma Agent – the top commercial solution! This demonstrates that Agentless can be complementary to existing commercial agent-based setups.
@@
Localization performance. In real-world software development, apart from directly fixing the issue, providing the correct edit location to human developers is extremely helpful for debugging. As such, we examine the locations of the patches generated by each technique compared with the ground truth patch. We note here that it is possible to fix a bug in a different location than the ground truth, however comparing against the ground truth patch can still serve as an approximate measure. Table 1 additionally shows the percentage of submitted patches with correct locations for each tool, across line, function, and file levels. We first observe that the percentage of patches with correct locations correlates heavily with the solve rate. Interestingly, the highest result in terms of file-level location is OpenCSG StarShip at 90.0%, significantly higher than even the best-performing approaches while at the same time having a relatively low solve rate (23.67%). As OpenCSG StarShip is a commercial product that does not provide source code or detailed trajectories, it is difficult to explain this huge difference between localization and repair performance. In terms of localization performance, by using our simple hierarchical approach, Agentless remains very competitive compared with previous agent-based approaches (best function-level, and second-best file- and line-level localization among all open-source approaches).
Table 2: Performance of different localization steps of Agentless.
Step Contains GT LoC Avg. \$
1. file level 77.7% 3,305 \$0.02
2. related classes + functions 55.3% 813 \$0.02
3. edit locations 50.8% 246 \$0.05
Ablation study on components of Agentless. Next, we look at how each component in both localization and repair phases contributed to the final Agentless performance. Table 2 shows the performance of each of the 3 step in Agentless's localization phase (for step-3, the metrics are averaged across two sets of locations, with the cost being the total cost). We show after each localization step the percentage of ground truth edit locations that still remains, the lines of code in each localization, and the average dollar cost of each step. We observe that Agentless is able to localize the ground truth file in 77.7% of cases; however, using all of the localized files leads to a huge number of code lines as part of the context. As such, in our second localization step, we localize to relevant classes and functions, and are able to drastically reduce the context window. Finally, Agentless localizes to the exact edit locations needed to achieve even more context reduction without losing much of the localization accuracy. Furthermore, we observe that by using hierarchical localization steps, Agentless can successfully minimize the cost while performing effective localization.
Table 3: Performance of different repair setups of Agentless.
Setup Performance Avg. \$
single sample 70 (23.33%) \$0.11
+ multiple samples & maj. voting 78 (26.00%) \$0.34
+ test filtering (full Agentless) 82 (27.33%) \$0.34

### v2

@@ 3.2 Repair [Uncaptioned image]
In the repair stage, the goal is to produce the correct patch to solve the issue. Following existing work on LLM-based program repair Xia and Zhang [2022]; Kolak et al. [2022]; Xia et al. [2023b]; Jiang et al. [2023], we first utilize the identified edit locations and construct a context window of code snippets to provide to the LLM for repair. For example, if the identified location was a class from line 40 to 78, we would produce a context window of [40 - x, 78 + x] where x denotes the context window size. The intuition behind adding the additional code before and after the identified location is to provide the LLM with relevant contextual information for better program repair Xia and Zhang [2022]. If multiple edit locations are identified, we would concatenate these context windows together separated with ``...'' to indicate missing context in the middle.
@@
Using the code snippets, we then ask the LLM to generate patches to solve the issue. However, instead of directly producing the entire code snippet to replace the entire given context, Agentless asks the LLM to generate a Search/Replace edit Gauthier [2024]: a simple diff format to efficiently create each patch. Figure 3 shows an example of the Search/Replace format containing two main parts: 1) search: the original code snippet we want to replace and 2) replace: the replacement code snippet we want to replace with. To apply the generated Search/Replace diff to the original file, we can simply match the search code snippet and replace it with the replacement. This simple diff format avoids generating the complete code and instead focuses on producing small edits, which are not only more cost-efficient, but also more reliable and accurate (less chances for hallucination). For each issue, Agentless uses the LLM to generate multiple potential patches (starting with greedy and then sample multiple patches with higher temperature).
3.3 Patch Validation [Uncaptioned image]
Figure 4: Example reproduction test.
3.3.1 Reproduction test generation.
Since Agentless generates multiple candidate patches per issue, we need a way to select a final patch for submission. Please note that under the realistic SWE-bench setup, the original project codebase can only provide regression tests and not any reproduction tests (i.e., bug-triggering tests). This is because the issue has just been raised and the developers have not added any additional tests to trigger the issue. As such, different from the Generate-and-Validate program repair setup Long and Rinard [2016], we do not have access to bug-triggering tests.
Following prior work Ruan et al. [2024]; Chen et al. [2024], Agentless generates additional reproduction test to help with patch selection. More specifically, Agentless leverages the LLM to synthesize a complete testing file that attempts to both reproduce the original issue described in the issue description, as well as verify whether the issue has been fixed. Figure 4 shows an example of the reproduction test that we want the model to synthesize. If the issue is reproduced, the test outcome should print Issue reproduced. On the other hand, the test should output Issue resolved if the issue has been fixed. We also include another output of Other issues if the test runs into any unexpected issues. To generate the reproduction test, Agentless provides the original issue description with an example reproduction test to demonstrate the test format. Similar to repair, we also sample multiple candidate reproduction tests and then execute each test on the original repository to filter any tests that do not output Issue reproduced. Finally, we normalize each test (remove comments, extra spaces, and normalize test names) and then select the test with the highest number of occurrence as the final reproduction test for each issue.
3.3.2 Patch selection.
Using the generated reproduction tests, we start our patch selection process to pick the final submission patch. Agentless first runs all the existing tests in the repository to identify a set of passing tests that successfully pass in the original codebase. However, not all of those passing tests should be considered as regression tests since solving the issue may require changing some of the existing functionalities. Therefore, Agentless provides the list of passing tests to the LLM and ask it to identify any tests that should not be ran to check if the issue has been correctly fixed (i.e., the tests that may be updated/patched during issue fixing). After removing the LLM-identified non-regression tests, we obtain a final set of regression tests. We then run the set of regression tests on all the generated patches. Agentless then keeps the patches with the lowest number of regression failures. For those patches, Agentless then runs the selected reproduction test and only keeps patches that output Issue resolved. Meanwhile, because the reproduction test is generated by the LLM and can potentially be incorrect/imprecise, it could be possible that no patch can pass the reproduction test; in that case, Agentless will fall back on only using the regression test results for selection. Agentless then applies a re-ranking approach using majority voting: We first normalize each patch to ignore surface-level differences (e.g., extra spaces, newlines, and comments), and then select the patch with the highest number of occurrences as the final patch for submission. More specifically, to standardize the patch, we begin by parsing both the old and new code (after applying the patch) into abstract syntax trees. Next, we unparse the trees into a canonical source code format with docstrings removed. Finally, we compute the textual diff between the standardized old and new code to get the normalized patch.
@@ 3.3 Patch Validation [Uncaptioned image]
Agentless solves repository-level issues using a simple step-by-step procedure. We note here that none of the techniques used by Agentless in isolation are revolutionary, but instead Agentless smartly combines existing techniques to construct an easy-to-understand approach. Different from prior autonomous agent-based tools that involve complex interactions with the environment, Agentless uses a simplistic three-phase approach to localize, repair, and validate without relying on any agents for decision-making. By conducting localization in a hierarchical manner, Agentless can efficiently and effectively compute the fine-grained locations for editing. Agentless then performs repair by sampling multiple patches using a simple diff format. Agentless's patch validation approach can further aid the patch selection process by producing reproduction tests that can help verify if the issue is fixed.
@@ 4 Experimental Setup
Datasets. We evaluate Agentless and baselines using the popular SWE-bench dataset Jimenez et al. [2024a] to test the ability to solve real-world software engineering issues. Each problem in SWE-bench requires submitting a patch to solve the underlying issue described in the input issue description. In particular, we focus on the widely used SWE-bench Lite version swe [2024], containing 300 self-contained problems with better quality. Furthermore, we also conduct a detailed study (Section 6.1) on the SWE-bench Lite benchmark to not only demonstrate potential issues and biases, but also produce a more rigorous filtered set of problems for better evaluation.
@@
Implementation. We implement Agentless using GPT-4o (gpt-4o-2024-05-13OpenAI [2024a]. By default, we query the LLM with greedy decoding. During sampling, we use a sampling temperature of $0.8$. For the embedding-based retrieval method, we implement our approach using LlamaIndex LlamaIndex [2024]. We use OpenAI's text-embedding-3-small OpenAI [2024b] model to compute the embedding with chunk size of 512 and chunk overlap of 0. For each issue, we first localize to the top three suspicious files, and then localize to an unrestricted number of suspicious classes and functions within these files, all using greedy decoding. Next, to maximize the chances of finding the correct edit locations, we draw four samples of edit locations per issue (i.e., the third step in the localization phase). This gives us 4 separate sets of edit locations per issue. For each set, we adopt a context window of $\pm$ 10 lines around each edit location, and generate 10 patches (1 greedy and 9 samples). This results in a total of 40 patches per bug. We adopt the same Search/Replace edit format from prior work Gauthier [2024], and use the built-in Python ast library ast [2024] to perform parsing in our normalization step. To generate the reproduction tests, we also generate 40 samples (1 greedy and 39 samples) in total prior to patch selection (described in Section 3.3.1). The regression tests are obtained by first running all the tests to obtain a set of passing tests that successfully pass in the original repository and then use the LLM to identify any non-regression tests (described in Section 3.3.2). We do not directly use the provided list of regression tests already identified in the PASS_TO_PASS field of SWE-bench as requested by the SWE-bench maintainers [^1] If we directly use the PASS_TO_PASS tests the performance on SWE-bench Lite will be 98.. We modify the official SWE-bench evaluation setup to be able to freely execute arbitrary regression and reproduction tests.
@@
Baselines. We compare Agentless against 26 agent-based approaches. These baseline tools represent the state-of-the-art performance on SWE-bench. We include state-of-the-art open-source as well as commercial or closed-source baselines (indicated via a [Uncaptioned image] ). We note here that the majority of the closed-source baselines do not provide any trajectories, just the submission patches. Therefore, we cannot verify the steps taken to arrive at the final patches. Moreover, we also include a simple agentless baseline using retrieval-augmented generation (RAG) proposed as part of SWE-bench Jimenez et al. [2024a] for comparison. In this case, the agentless baseline uses the LLM to directly generate a patch file by providing it with the file content of the most relevant files, retrieved using BM25 Robertson et al. [2009].
@@
Metrics. Following prior work Zhang et al. [2024c], we report 1) % Resolved: the percentage of resolved problems in the benchmark, 2) Avg. \$ Cost: average inference cost of running the tool, and 3) Avg. # Tokens: average number of input and output tokens used to query to LLM. Additionally, we also report the % Correct Location: the percent of problems where the patch produced by the tool covers the edit location(s) of the ground truth developer patch. We compute this metric over three granularities: file, function, and line. We report that a patch contains the correct location if it edits a superset of all locations in the ground truth patch. For baseline tools, we directly use the reported results either from the official leaderboard Jimenez et al. [2024b] or from the tool's official paper/repository.
@@ 5 Evaluation
Table 1: Results on SWE-bench Lite. Note: [Uncaptioned image] indicates approaches that are closed-source (i.e., source code is not released). [Uncaptioned image] indicates approaches built on top of an earlier version of our Agentless. '-' indicates that the relevant information to compute this has not been released. [Uncaptioned image] indicates that multiple models are used, but some of them are not specified. Claude 3.5 S is short for Claude 3.5 Sonnet.
Tool LLM % Resolved
Avg.
\$ Cost
Avg.
# Tokens
% Correct Location
@@
CodeStory Aide cod [2024] [Uncaptioned image] [Uncaptioned image] GPT-4o+ [Uncaptioned image] Claude 3.5 S 129 (43.00%) - - 41.7% 58.7% 72.0%
Bytedance MarsCode Liu et al. [2024b] [Uncaptioned image] NA 118 (39.33%) - - 42.7% 58.0% 79.7%
Honeycomb hon [2024] [Uncaptioned image] NA 115 (38.33%) - - 44.3% 57.0% 69.3%
MentatBot men [2024] [Uncaptioned image] [Uncaptioned image] GPT-4o 114 (38.00%) - - 37.3% 53.3% 69.3%
Gru gru [2024] [Uncaptioned image] NA 107 (35.67%) - - 38.3% 54.3% 75.0%
Isoform iso [2024] [Uncaptioned image] NA 105 (35.00%) - 41,963 38.7% 55.3% 72.0%
SuperCoder2.0 sup [2024] [Uncaptioned image] NA 102 (34.00%) - - 41.7% 63.7% 65.7%
Alibaba Lingma Agent lin [2024] [Uncaptioned image]
[Uncaptioned image] GPT-4o+
Claude 3.5 S
99 (33.00%) - - 40.0% 58.7% 75.0%
@@
Amazon Q Developer-v2 ama [2024] [Uncaptioned image] NA 89 (29.67%) - - 40.3% 52.0% 74.3%
SpecRover Ruan et al. [2024] [Uncaptioned image] [Uncaptioned image] GPT-4o+ [Uncaptioned image] Claude 3.5 S 93 (31.00%) \$0.65 - - - -
@@
MASAI Arora et al. [2024] [Uncaptioned image] NA 84 (28.00%) - - 38.7% 56.3% 75.0%
SIMA sim [2024] [Uncaptioned image] [Uncaptioned image] GPT-4o 83 (27.67%) \$0.82 - 37.0% 54.0% 79.0%
@@
AutoCodeRover-v2 aut [2024] [Uncaptioned image] GPT-4o 92 (30.67%) - - 35.0% 52.3% 69.3%
RepoGraph rep [2024] [Uncaptioned image] GPT-4o 89 (29.67%) - - 36.7% 51.3% 71.0%
Moatless moa [2024]
Claude 3.5 S
80 (26.67%) \$0.17 - 38.7% 54.7% 78.7%
[Uncaptioned image] GPT-4o 74 (24.67%) \$0.14 - 36.0% 52.0% 73.0%
OpenDevin+CodeAct v1.8 ope [2024b]
Claude 3.5 S
80 (26.67%) \$1.14 - 38.0% 49.7% 67.3%
Aider Gauthier [2024] [Uncaptioned image] GPT-4o+ [Uncaptioned image] Claude 3.5 S 79 (26.33%) - - 35.3% 50.0% 69.7%
SWE-agent Yang et al. [2024a]
Claude 3.5 S
69 (23.00%) \$1.62 521,208 40.7% 54.3% 72.0%
[Uncaptioned image] GPT-4o 55 (18.33%) \$2.53 498,346 29.3% 42.3% 58.3%
@@
AppMap Navie app [2024] [Uncaptioned image] GPT-4o 65 (21.67%) - - 29.7% 44.7% 59.7%
AutoCodeRover Zhang et al. [2024c] [Uncaptioned image] GPT-4 57 (19.00%) \$0.45 38,663 29.0% 42.3% 62.3%
RAG Yang et al. [2024a]
Claude 3 Opus
@@
[Uncaptioned image] GPT-3.5 1 (0.33%) - - 6.3% 11.3% 27.3%
Agentless [Uncaptioned image] GPT-4o 96 (32.00%) \$0.70 78,166 35.3% 52.0% 69.7%
5.1 Performance on SWE-bench Lite
@@
Table 1 shows the main evaluation result of Agentless and prior agent-based approaches on SWE-bench Lite. We observe that Agentless is able to solve 96 out of 300 problems (32.00%). While this is not the highest percentage of problems solved on SWE-bench Lite, Agentless is extremely competitive compared with prior agent-based approaches while using a much simpler design and overall technique. It is important to note here that many of the top techniques are closed-source/commercial and did not release any source code to reproduce experiments or even trajectories for further verification. Compared with all open-source approaches, Agentless is able to achieve the highest performance of 32.00% (96 / 300) on SWE-bench Lite. Additionally, Agentless only costs on average \$0.70, which is less than most prior agent-based approaches. Comparing against the RAG agentless baselines, we see that while Agentless costs slightly more, Agentless is also able to fix way more issues.
Figure 5: Venn diagram for issue fixes.
5.1.1 Unique issues fixed.
@@
Figure 5 shows the unique issues solved by Agentless compared with the top-performing closed-source / commercial and open-source approaches (``Others'' indicates all other approaches within each category). First, we see that compared to the open-source agent-based tools, Agentless is able to fix 2 issues that no other tools can resolve, showing the success of using a simple agentless approach in solving difficult issues. Furthermore, even when compared with high-performing commercial approaches, Agentless is still able to offer a unique fix. This low number of unique fixes can be attributed to the fact there are already tools built on top of earlier versions of Agentless (e.g., Isoform iso [2024]) or partly inspired by Agentless (e.g., Bytedance MarsCode Liu et al. [2024b]), thereby reducing the unique issues resolved by Agentless. Nevertheless, our results show that Agentless can be competitive/complementary compared to existing agents.
5.1.2 Localization performance.
@@
In real-world software development, apart from directly fixing the issue, providing the correct edit location to developers is extremely helpful for debugging. As such, we examine the locations of the patches generated by each technique compared with the ground truth patch. We note here that it is possible to fix a bug in a different location than the ground truth, however comparing against the ground truth patch can still serve as an approximate measure. Table 1 additionally shows the percentage of submitted patches with correct locations for each tool, across line, function, and file levels. We first observe that the percentage of patches with correct locations correlates heavily with the solve rate. Interestingly, the highest result for file-level location is OpenCSG StarShip at 90.0%, significantly higher than even the best-performing approaches while at the same time having a relatively low solve rate (23.67%). As OpenCSG StarShip is a commercial product that does not provide source code or detailed trajectories, it is difficult to explain this huge difference between localization and repair performance. In terms of localization performance, by using our simple hierarchical approach, Agentless remains very competitive compared with previous agent-based approaches.
5.1.3 Reproduction test results.
Agentless uses regression and generated reproduction tests to perform filtering in order to select the final submission patch. Therefore, we evaluate the quality of our generated reproduction tests. We note here that as described in Section 3.3.1, we only use the generated reproduction test if it can successfully reproduce the original issue in the original repository. Out of the 300 problems in SWE-bench Lite, Agentless is able to produce 213 reproduction tests that output the required reproduction message when evaluated on the original repository. However, these tests might still be incorrect as a correct reproduction should also be able to verify that the issue has been correctly resolved. To evaluate this, we directly apply the ground truth patch provided in SWE-bench  [^2] Note that the ground truth patch is only applied here to evaluate the test quality, and is not used during the Agentless process for selecting the reproduction tests.. We found that only 94 tests correctly output the Issues resolved message after applying the ground truth patches. This steep drop-off can be partially explained as sometimes the issue description provided in the problem may not contain enough information to generate complete test cases for validating a correct solution. However, this problem is partially mitigated as Agentless takes a conservative approach of requiring all patches to pass the regression test suite first and will remove a patch if it cannot pass the regression tests but passes on the generated reproduction test. This reduces the likelihood of an incorrect reproduction test selecting a correct patch. Furthermore, if all generated patches cannot pass the reproduction test, Agentless will fall back on only using the regression test results for selection. In Section 5.2.3 we closely examine the impact of using both regression and reproduction test for patch selection has on the performance.
Table 2: Performance of different localization steps.
Method Contains GT LoC Avg. \$
File level localization
Prompting-based 78.67% 3,221 \$0.02
Embedding-based
(w/o irrelevant filtering)
67.67% 3,388 \$0.05
Embedding-based (w/ irrelevant filtering) 70.33% 3,622 \$0.04
Combined 81.67% 3,424 \$0.06
Related element localization
Complete file 53.67% 778 \$0.15
Skeleton format 58.33% 698 \$0.02
Edit location localization
Greedy 50.67% 189 \$0.06
Direct from file-level 47.00% 208 \$0.18
Multi-samples merged 56.33% 342 \$0.07
Multi-samples 49.67% 165 \$0.07
49.33% 180
49.33% 168
48.33% 213
5.1.4 Adoption of Agentless.
Although released very recently, Agentless has already received widespread adoption. In Table 1, there are two baseline tools (indicated via [Uncaptioned image] ) already built upon earlier versions of our Agentless: RepoGraph rep [2024] is an open-source tool combining Agentless with repository-level graph, while Isoform iso [2024] is a closed-source commercial tool also built upon Agentless. Additionally, Bytedance MarsCode Liu et al. [2024b] is partly inspired by Agentless in patch sampling/selection and OpenDevin ope [2024b] is in the process of integrating Agentless into their ecosystem. Furthermore, Agentless has also been adopted by OpenAI as the default approach when showcasing the real-world coding performance of GPT-4o on their new SWE-bench Verified benchmark Chowdhury et al. [2024], where Agentless also achieves the best performance compared with all other studied agent-based solutions. At the time of finishing up this draft, OpenAI just released the new o1 model family and also adopted Agentless as the top/default approach to showcase their performance on SWE-bench OpenAI [2024c].
5.2 Ablation study on components of Agentless
Next, we look at how each component in the localization, repair, and patch validation phases contributed to the final Agentless performance. Unless otherwise specified, we vary the configuration of one component while using the default parameters for all other settings.
5.2.1 Localization ablation.
Table 2 shows the performance and cost for each of the 3 steps in Agentless's localization phase. We show after each localization step the percentage of problems whose ground truth edit locations remain in the location set (``Contains GT''), the average lines of code of each location set (``LoC''), and the average dollar cost of each step (``Avg.\$''). The bold method indicates the default setting of Agentless. First, we examine the different configurations of file level localization. To start with, for the retrieval method based on embeddings, we see that without including the irrelevant folder filtering to remove irrelevant folders for embedding (described in Section 3.1.1), both the performance and cost become worse. This demonstrates the importance of limiting the number of files to consider during embedding and focusing on essential parts of the repository for more cost-efficient and effective localization. We see that using the prompting-based or the embedding-based retrieval method alone can locate the ground truth file in 78.7% and 67.7% of cases respectively. This can be further improved by combining them to obtain 81.7% correct file localization, showing that prompting-based and embedding-based retrieval methods can complement each other in identifying different sets of relevant files.
Table 3: Performance of different repair setups.
Method Performance Avg. \$
Greedy location (40 samples) 88 (29.33%) \$0.22
Multi-samples merged
(40 samples)
85 (28.33%) \$0.24
Multi-samples (4 x 10 samples) 96 (32.00%) \$0.29
Using all of the localized files leads to a large context window (>>3000). As such, in our second localization step, we localize to the relevant classes and functions, and are able to drastically reduce the context window (<<800). We compare our input of using skeleton format (described in Section 3.1.2) to provide a more concise representation with the baseline of using the complete file content. We observe that by using the complete file content, not only is the cost much higher but also the number of localized groundtruth issues is reduced. The reason is most likely that LLMs cannot handle long context very well, so providing the entire file contents can confuse the model. Conversely, by using a more concise representation as the input, we can effectively localize the correct related locations that are needed for inspection and editing.
Next, Agentless localizes to the exact edit locations needed to achieve even more context reduction without losing much of the localization accuracy. We compare the different ways we can perform the edit location localization: 1) ``Greedy'': using greedy decoding to obtain one set of edit locations, 2) ``Direct from file-level'': directly go from file-level localization to the edit locations (instead the default of localizing from file-level to related elements and then to edit locations), 3) ``Multi-samples merged'': sample multiple sets of edit locations and merging them into one set, and 4) ``Multi-samples'': sample multiple sets of edit locations. We first observe that by directly going from file-level to the edit locations, both the cost and performance are worse. The reason is that the model can become confused when providing a large context, demonstrating the importance of our hierarchical localization design. We also find that when merging multiple samples together, the amount of ground truth localized is higher but at the expense of having to add more context as the input during the repair phase. Our default settings also sample the edit locations multiple times, however instead of merging, we perform repair on them separately to make use of the fact each sampled location set can provide similar localization performance while also limiting the input context. In short, by using hierarchical localization steps, Agentless can successfully minimize the cost while performing effective localization.
5.2.2 Repair ablation.
We now look at the impact of our different repair setups on the final performance. Table 3 shows the different settings and inputs for our repair phase with their performance and cost.Starting with using the greedy location set generated in the edit location generation stage, we observe that we can already achieve very high performance of more than 88 issues fixed. Similarly, for the ``Multi-samples merged'' where we merged multiple location sets into one, we can also achieve comparable performance. The performance can be further improved by considering each sampled locations separately (to generate 10 candidate patches each) when performing repair to achieve 96 fixes. The reason is that each different location sets may localize different ground truth locations and provide different context that can be helpful to fix specific issues. By using different edit locations and combining with our extensive test filtering and selection stage, Agentless can drastically improve the repair performance.
Figure 6: Repair performance as the number of patch samples increases.